OpenStudy (anonymous):

tan10 + tan35 / 1- tan10tan35

OpenStudy (anonymous):

this is probably either the addition angle formula, or the subtraction i can never keep them straight, look it up my guess is the addition, so it is probably \[\tan(45^\circ)\] but i would not swear to it

OpenStudy (anonymous):

So can you walk me thru the steps or no?

OpenStudy (anonymous):

Help plz

OpenStudy (anonymous):

Trigonometry Identities I believe this is

OpenStudy (perl):

$$ \large\tan ( A + B ) = \frac{\tan A + \tan B }{1 - \tan A \cdot \tan B } $$

OpenStudy (anonymous):

Which wud b A and B

OpenStudy (perl):

$$ \large\tan ( A + B ) = \frac{\tan A + \tan B }{1 - \tan A \cdot \tan B } \\~\\ \large\tan ( 10 + 35 ) = \frac{\tan 10 + \tan 35 }{1 - \tan 10 \cdot \tan 35 } $$

OpenStudy (perl):

but tan(10 + 35) = tan(45) = 1

OpenStudy (anonymous):

so then tan 45 = tan 45/1- tan 45

OpenStudy (perl):

not quite. tan 45 = tan(10 + 35) = ( tan 10 + tan 35) / ( 1- tan 10 * tan 35) the point here is that tan 45 is a well known angle , so you're basically done at that point. tan 45 is 1

OpenStudy (perl):

this is a reverse logic problem

OpenStudy (perl):

you wouldn't use this formula to find tan 45, because it makes finding tan 45 more complicated

OpenStudy (anonymous):

so the answer is tan=1

OpenStudy (perl):

$$ \Large \frac{\tan 10 + \tan 35 }{1 - \tan 10 \cdot \tan 35 }= \tan ( 10 + 35 ) = \tan(45) = 1 $$

OpenStudy (anonymous):

can u help with another prob??

OpenStudy (perl):

ok

OpenStudy (anonymous):

want me to create a new question n tag u or just post it here?

OpenStudy (perl):

sure

OpenStudy (perl):

new question

OpenStudy (anonymous):

I just posted the new question and tagged u

OpenStudy (perl):

I don't see the tag, for some reason.

OpenStudy (anonymous):

try looking on the open question feed